# Breadcrumbs Reasoning: Memory-Efficient Reasoning with Compression Beacons

### v1

@@ 5 Results
We compare the BR policies across different compression ratios to the $\pi_{\rm RL}$ that we use as the distillation source. We maintain the same source policy for all our experiments to make this comparison valid. Table 2 reports the results for a fixed maximum generation length (i.e., time steps) of 1,000.
While on Countdown and StarGraph all compression ratios work well, performance on LinSys varies greatly across compression ratios. As expected, higher compression ratios lead to worse performance. For LinSys, BR lags behind the teacher for both models. We speculate this is due to differences in the reasoning challenge compared to Countdown and StarGraph. Overall, given a fixed response length, BR preserves performance across tasks by 67.1–94.0% for Qwen and 65.1–84.5% for Phi while using only 2–32x fewer KV cache entries at generation time.
@@
The training-free methods TOVA and StreamingLLM consistently underperform across tasks. On Countdown, their performance drops dramatically with higher compression (e.g., TOVA falls from 0.574 at 2× to 0.172 at 8× for Qwen; StreamingLLM remains below 0.32 across all settings). On StarGraph, the gap between the baselines and out approach is even more severe, with StreamingLLM dropping below 0.1 accuracy in nearly every configuration. In LinSys, both baselines fail to reach even a single correct output. These results highlight the limitation of training-free cache eviction methods: for tasks requiring long coherent reasoning chains, simply truncating past tokens eliminates essential intermediate steps, from which the model is unable to recover.
@@
We compare two strategies for distilling from $\pi_{RL}$ in Figure 5. In our joint RL-distillation training, BR is distilled online from the rollouts of the teacher $\pi_{\rm RL}$ as it learns; in late distillation, compression policies are trained only on trajectories sampled from a final checkpoint of $\pi_{\rm RL}$. BR achieves equivalent or superior performance in 26 of the 30 configurations tested, and very close performance on the other four. This demonstrates that it effectively piggybacks on the same data used to train $\pi_{\rm RL}$. This eliminates the need for additional distillation data, minimizes training overhead, and does not impose the need to decide a priori a number of training steps for $\pi_{\rm RL}$. We hypothesize that the superior performance may derive from the distributional shift of $\pi_{\rm RL}$ during RL training.
@@ 6 Discussion
Although at a fixed generation length our method shows a small performance loss, we demonstrate that when generating many more tokens under the same memory budget, performance often surpasses that of a non-compressing teacher. This is enabled by effective test-time scaling: when matching the KV-cache size, we are able to generate significantly more tokens. To some extent, these results show that we trade memory for time. Such trade-offs are common in efficiency-oriented methods — for example, speculative decoding (Leviathan et al., 2023; Chen et al., 2023) reduces latency but requires more memory, since multiple models must be loaded simultaneously.
There are several directions for future work, entailed from several areas where our approach can be improved. Foremost, our compression policy is not dynamic — one cannot adjust its compression rate as desired. It is advantageous to have a single model that supports different compression rates, and therefore can be used flexibly in different settings. This is an important direction for future work, and one that is relatively understudied in the compression literature. Our work also charts the direction for future work to improve test-time scaling and compression tradeoffs. Ideally, one can compress aggressively without needing to increase the number of reasoning steps. Our work exposes this tradeoff in reasoning models, and outlines the methodology to analyze it. Finally, as the space of accessible benchmarking scenarios in reasoning research develops, it is important to understand the behavior of our approach across a broader set of domains.
@@ Acknowledgments
We thank Nathan Godey for his insightful discussions and feedback, and Alessio Devoto for his helpful discussions regarding KV compression implementations (and for his work on the kvpress library). We thank the Cornell NLP group for helpful comments. This research was supported by NSF under grants No. 1750499 and OAC-2311521, NASA under award No. 20-OSTFL20-0053, the Academic Grant Program, the National Artificial Intelligence Research Resource (NAIRR) Pilot, the Frontera supercomputer supported by the National Science Foundation (award NSF-OAC 1818253) at the Texas Advanced Computing Center (TACC) at The University of Texas at Austin, and the Delta advanced computing and data resource which is supported by the National Science Foundation (award NSF-OAC 2005572). We gratefully acknowledge use of the research computing resources of the Empire AI Consortium, Inc, with support from the State of New York, the Simons Foundation, and the Secunda Family Foundation (Bloom et al., 2025). KB acknowledges this work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence.
@@ Appendix A Implementation Details
We use VeRL (Sheng et al., 2024) as the backbone for our training code. We adapt kvpress (Jegou et al., 2024), a recent library for Transformers KV cache compression for inference (e.g., for evaluation). We decouple RL and distillation during our experiments, so that the teacher data is the same for all BR policies and results can be compared more fairly. Otherwise, different experimental results may depend on the different quality of $\pi_{\rm RL}$ training, rather than on the methods being tested. We save all $\pi_{\rm RL}$ sampled trajectories to decouple training, and the top-100 logits for each time step, because saving all logits is extremely storage-intensive. In all training setups, this leads to consistently distilling more than 90% of the probability mass for each token, although it frequently reaches even higher levels throughout training.
@@ Appendix B Task Details
  • StarGraph: We use star graphs (Bachmann & Nagarajan, 2024) with branches of length 5 and up to 25 branches per graph. When creating the dataset, the number of branches for each graph is sampled uniformly from the range $[2,25]$. We empirically found that this variable complexity helps the reinforcement learning policy, $\pi_{\text{RL}}$, by allowing it to gradually learn to solve more complex graphs.
@@ C.1 $\pi_{\rm RL}$ Training
Table 3 provides the hyperparameters for PPO Schulman et al. (2017) and AdamW (Loshchilov & Hutter, 2019). The critic has the same architecture and initial weights as $\pi_{\rm RL}$.
@@ C.2 $\pi_{\rm BR}$ Training
Appendix D LLM Usage

### v3

@@ 5 Results
We compare the BR policies across different compression ratios to the $\pi_{\rm RL}$ that we use as the distillation source. We maintain the same source policy for all our experiments to make this comparison valid. Section 5 reports the results for a fixed maximum generation length (i.e., time steps) of 1,000.
While on Countdown and StarGraph all compression ratios work well, performance on LinSys varies greatly across compression ratios. As expected, higher compression ratios lead to worse performance. For LinSys, BR lags behind the teacher for both models. We speculate this is due to differences in the reasoning challenge compared to Countdown and StarGraph. Overall, given a fixed response length, SR BR preserves performance across tasks by 67.1–94.0% for Qwen and 65.1–84.5% for Phi while using only 2–32x fewer KV cache entries at generation time.
Multi-Ratio Training
We analyze the performance of MR BR, where a single model is trained to handle all compression ratios simultaneously. One might expect that learning a policy capable of varying compression granularities would be a harder optimization problem, potentially leading to lower performance compared to the specialized SR BR models. However, our results indicate the opposite. MR BR not only retains the flexibility to operate at any ratio at inference time but also outperforms SR BR on average. For example, on the Qwen model in Section 5, MR BR achieves an average accuracy of 69.6% across all tasks and compression ratios, compared to 68.1% for SR BR. This suggests that the joint training facilitates information sharing, enabling the model to transfer effective compression strategies across varying ratios.
@@
The training-free methods TOVA and StreamingLLM consistently underperform across tasks. On Countdown, their performance drops dramatically with higher compression (e.g., TOVA falls from 0.574 at 2× to 0.172 at 8× for Qwen; StreamingLLM remains below 0.32 across all settings). On StarGraph, the gap between the baselines and out approach is even more severe, with StreamingLLM dropping below 0.1 accuracy in nearly every configuration. In LinSys, both StreamingLLM and TOVA fail to reach even a single correct output, while SnapKV and PyramidKV are still below 2% accuracy even with the lowest compression ratio. These results highlight the limitation of training-free cache eviction methods: for tasks requiring long coherent reasoning chains, simply truncating past tokens eliminates essential intermediate steps, from which the model is unable to recover.
Figure 4: Failure mode analysis on LinSys. We classify error types of Qwen with SR BR across increasing compression ratios. The results indicate that the performance drop in LinSys is primarily driven by arithmetic errors, which increase significantly with compression, rather than a failure to memorize or retrieve information.
Why Does BR Struggle with LinSys?
While BR retains most of the accuracy in Countdown and StarGraph, LinSys plateaus earlier, especially for larger compression ratios. The main difference between this task and the other two is that it requires careful arithmetic and algebraic manipulation. However, it could also be that beacons forget more in this task than in the others. We test the two possibilities (BR struggles with arithmetic vs. beacons fail to memorize) with an LLM-as-a-Judge (NEURIPS2023_91f18a12) pipeline (detailed in Appendix D) and report our findings in Figure 4. Interestingly, we find that the vast majority of mistakes are attributed to arithmetic mistakes, and their number increases significantly with increased compression. One potential factor is that compressing numbers might be harder, and also that our compression method might break some arithmetic circuits of the model, and that our training is not long enough to alleviate these issues.
@@
We compare two strategies for distilling from $\pi_{RL}$ in Figure 5 (for SR BR). In our joint RL-distillation training, BR is distilled online from the rollouts of the teacher $\pi_{\rm RL}$ as it learns; in late distillation, compression policies are trained only on trajectories sampled from a final checkpoint of $\pi_{\rm RL}$. BR achieves equivalent or superior performance in 26 of the 30 configurations tested, and very close performance on the other four. This demonstrates that it effectively piggybacks on the same data used to train $\pi_{\rm RL}$. This eliminates the need for additional distillation data, minimizes training overhead, and does not impose the need to decide a priori a number of training steps for $\pi_{\rm RL}$. We hypothesize that the superior performance may derive from the distributional shift of $\pi_{\rm RL}$ during RL training.
Qwen Phi
Method 2 4 8 16 32 2 4 8 16 32
[…]
Sr Br (Ours) 1.990 1.546 1.593 1.646 1.468 2.029 1.486 1.518 1.506 1.164
Mr Br (Ours) 1.999 1.665 1.507 1.396 1.366 1.821 1.574 1.307 1.306 1.220
[…]
PyramidKV 1.844 2.212 1.858 1.495 1.649 1.823 4.817 10.341 7.669 10.687
SnapKV 2.607 2.364 1.748 1.664 1.504 1.738 4.347 14.438 10.575 15.234
TOVA 1.880 2.353 2.751 2.563 2.470 1.655 2.184 3.188 3.850 3.890
StreamingLLM 2.326 3.210 4.988 5.281 6.349 2.191 3.815 7.354 3.485 3.857
Table 2: Latency increase across tasks. The values represent the relative slowdown compared to the Teacher model (lower is better), given a maximum generation KV cache size of 1,000 entries (i.e., up to 32x tokens more with a ratio of 32). Columns represent the compression ratio. Bold indicates the best result; underlined indicates the second best.
@@ 6 Discussion
Although at a fixed generation length our method shows a small performance loss, we demonstrate that when generating many more tokens under the same memory budget, performance often surpasses that of a non-compressing teacher. This is enabled by effective test-time scaling: when matching the KV-cache size, we are able to generate significantly more tokens. To some extent, these results show that we trade memory for time. Such trade-offs are common in efficiency-oriented methods — for example, speculative decoding (leviathan2023speculative; chen2023accelerating) reduces latency but requires more memory, since multiple models must be loaded simultaneously.
There are several directions for future work, entailed from several areas where our approach can be improved. While Multi Ratio Breadcrumbs Reasoning supports various compression rates, it lacks dynamic adaptivity. It does not automatically select the most appropriate ratio, and once a ratio is chosen, it remains fixed for the duration of the generation rather than adjusting dynamically across the sequence. This is an important direction for future work, and one that is relatively understudied in the compression literature. Our work also charts the direction for future work to improve test-time scaling and compression tradeoffs. Ideally, one can compress aggressively without needing to increase the number of reasoning steps. Our work exposes this tradeoff in reasoning models, and outlines the methodology to analyze it. Another promising direction is combining our method with the orthogonal CoT shortening approach (aggarwal2025l; kang2025c3ot; ma2025cot; shen2025dast; yan2025inftythink; munkhbat2025self; xia2025tokenskip), leading to savings both from shortening and compression, while still retaining the ability to for expressive reasoning through length merrill2023expressive. Finally, as the space of accessible benchmarking scenarios in reasoning research develops, it is important to understand the behavior of our approach across a broader set of domains.
@@ Acknowledgments
We thank Nathan Godey for his insightful discussions and feedback, and Alessio Devoto for his helpful discussions regarding KV compression implementations (and for his work on the kvpress library). We thank the Cornell NLP group for helpful comments. This research was supported by a gift to the LinkedIn–Cornell Bowers Strategic Partnership, NSF under grants No. 1750499 and OAC-2311521, NASA under award No. 20-OSTFL20-0053, the Academic Grant Program, the National Artificial Intelligence Research Resource (NAIRR) Pilot, the Frontera supercomputer supported by the National Science Foundation (award NSF-OAC 1818253) at the Texas Advanced Computing Center (TACC) at The University of Texas at Austin, and the Delta advanced computing and data resource which is supported by the National Science Foundation (award NSF-OAC 2005572). We gratefully acknowledge use of the research computing resources of the Empire AI Consortium, Inc, with support from the State of New York, the Simons Foundation, and the Secunda Family Foundation (Bloom2025EmpireAI). KB acknowledges this work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence.
@@ Appendix A Implementation Details
We use VeRL (sheng2024hybridflow) as the backbone for our training code. We adapt kvpress (devoto2025expectedattention), a recent library for Transformers KV cache compression for inference (e.g., for evaluation). We decouple RL and distillation during our experiments, so that the teacher data is the same for all BR policies and results can be compared more fairly. Otherwise, different experimental results may depend on the different quality of $\pi_{\rm RL}$ training, rather than on the methods being tested. We save all $\pi_{\rm RL}$ sampled trajectories to decouple training, and the top-100 logits for each time step, because saving all logits is extremely storage-intensive. In all training setups, this leads to consistently distilling more than 90% of the probability mass for each token, although it frequently reaches even higher levels throughout training.
Overview of Beacon Learning
Before training, we add to the vocabulary of the model a new token, the beacon. In the embedding matrix of the model, the beacon token is initialized as the average of all other embeddings. The model learns to compress normal tokens’ key/value activations because, thanks to the attention mask in Figure 2, otherwise those activations’ information would not be accessible from future tokens.
@@ Appendix B Task Details
  • StarGraph: We use star graphs (bachmann2024pitfalls) with branches of length 5 and up to 25 branches per graph. When creating the dataset, the number of branches for each graph is sampled uniformly from the range $[2,25]$. We empirically found that this variable complexity helps the reinforcement learning policy, $\pi_{\text{RL}}$, by allowing it to gradually learn to solve more complex graphs.
@@ C.1 $\pi_{\rm RL}$ Training
Table 3 provides the hyperparameters for PPO schulman2017proximalpolicyoptimizationalgorithms and AdamW (loshchilov2019decoupledweightdecayregularization). The critic has the same architecture and initial weights as $\pi_{\rm RL}$.
@@ C.2 $\pi_{\rm BR}$ Training
Appendix D Analysis
LLM-as-a-Judge Pipeline
We analyze failures of Qwen on the LinSys task in two stages. First, for every incorrect generation we pair the model’s reasoning trace with the ground truth and prompt a verifier LLM (Qwen3 30B A3B Thinking 2507) to identify the first erroneous step; the resulting snippet (e.g., “computed 1003100100-3100 as 2999-2999”) is stored. Second, we pass each extracted snippet to a classification judge that sees only the mistake text plus a fixed taxonomy (Arithmetic, Equation Misreading & Fabrication, Algebraic Manipulation, Output, Other), that we identify by manual inspection of the mistakes (and add Other to allow for cases that we do not explicitly categorize). Using best-of-$n$ sampling (n=3n{=}3), the judge assigns the most specific category.
Appendix E Additional Results
We report in Appendix E the accuracy area under the curve (AUAC) metric with varying the maximum cache size up to 1,000.
Figure 6 shows the Pareto curves for Single Ratio Breadcrumbs Reasoning.
Figure 6: Accuracy vs. KV Cache Size Single Ratio Breadcrumbs Reasoning retains most of the teacher’s performance while using significantly fewer KV cache entries, even outperforming the teacher when setting a fixed maximum KV cache size in Countdown and StarGraph.
Figure 7 illustrates the increase in accuracy of BR when increasing the generation length from 1,000 to 1,000 $\times c$ (i.e., KV cache size of 1,000 entries).
Qwen
Countdown LinSys StarGraph
Compr. Ratio 1 2 4 8 16 32 1 2 4 8 16 32 1 2 4 8 16 32
[…]
Teacher 0.377 - - - - - 0.276 - - - - - 0.513 - - - - -
[…]
Sr Br (Ours) - 0.465 0.533 0.555 0.539 0.513 - 0.451 0.532 0.368 0.347 0.286 - 0.726 0.845 0.906 0.918 0.926
Mr Br (Ours) - 0.446 0.515 0.576 0.576 0.560 - 0.439 0.507 0.424 0.391 0.316 - 0.713 0.839 0.912 0.921 0.935
[…]
PyramidKV - 0.375 0.201 0.111 0.077 0.011 - 0.000 0.000 0.000 0.000 0.000 - 0.487 0.459 0.491 0.463 0.526
SnapKV - 0.376 0.210 0.138 0.100 0.079 - 0.003 0.000 0.000 0.000 0.000 - 0.534 0.478 0.485 0.452 0.489
TOVA - 0.447 0.272 0.167 0.183 0.199 - 0.000 0.000 0.000 0.000 0.000 - 0.526 0.417 0.424 0.414 0.423
StreamingLLM - 0.007 0.015 0.024 0.049 0.090 - 0.000 0.000 0.000 0.000 0.000 - 0.038 0.032 0.023 0.041 0.105
Phi
Countdown LinSys StarGraph
Compr. Ratio 1 2 4 8 16 32 1 2 4 8 16 32 1 2 4 8 16 32
[…]
Teacher 0.433 - - - - - 0.142 - - - - - 0.498 - - - - -
[…]
Sr Br (Ours) - 0.506 0.557 0.581 0.604 0.597 - 0.373 0.421 0.322 0.205 0.187 - 0.636 0.739 0.786 0.785 0.792
Mr Br (Ours) - 0.480 0.526 0.582 0.568 0.609 - 0.375 0.454 0.408 0.278 0.258 - 0.631 0.710 0.765 0.732 0.738
[…]
PyramidKV - 0.417 0.225 0.180 0.019 0.000 - 0.011 0.000 0.000 0.000 0.000 - 0.503 0.366 0.292 0.327 0.358
SnapKV - 0.406 0.222 0.149 0.062 0.012 - 0.006 0.000 0.000 0.000 0.000 - 0.516 0.424 0.329 0.331 0.304
TOVA - 0.196 0.064 0.050 0.046 0.094 - 0.000 0.000 0.000 0.000 0.000 - 0.621 0.504 0.433 0.436 0.378
StreamingLLM - 0.014 0.018 0.060 0.150 0.300 - 0.000 0.000 0.000 0.000 0.000 - 0.137 0.017 0.029 0.089 0.097
(a) Model performance (AUAC) on long-context reasoning tasks. We report AUAC given a maximum cache size of 1,000 entries. Higher is better.
(a) Single Ratio Breadcrumb Reasoning
(b) Multi Ratio Breadcrumbs Reasoning
Appendix F LLM Usage